OpenStudy (anonymous):

Limit as x approaches 0 (3-3cosx)/(4x)

OpenStudy (anonymous):

lim x->0 3(1-cosx)/4x = 3/4 lim x-> 0 2 sin ^2 (x/2) /2(x/2) = 3/4 lim x-> 0 sin ^2 (x/2) /(x/2) = 3/4 x 1 =3/4

OpenStudy (anonymous):